OpenStudy (anonymous):

Two forces with magnitudes of 150 and 100 pounds act on an object at angles of 40° and 170° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer. (check my answer)

OpenStudy (anonymous):

First vector: 150<cos40,sin40> Second vector: 100<cos170,sin170> 150<cos40,sin40>+100<cos170,sin170>=150<0.77,0.64>+100<-0.98,0.17>=<115.5,96>+<-98,17>=<17.5,113> (resultant force's vector) ||<17.5,113>||=sqrt(17.5^2+113^2)=sqrt(306.25+12769)=sqrt(13075.25)=114.34 lbs (resultant force's magnitude)

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

I got magnitude 114.96 and direction 81.79 degrees

OpenStudy (anonymous):

How do you convert a vector into an angle?

OpenStudy (mertsj):

Tangent function

OpenStudy (anonymous):

Can you show me the steps please?

OpenStudy (mertsj):

For the resultant I got (16.426,113.783)

OpenStudy (mertsj):